OpenStudy (anonymous):

what are the perpendicular & parallel equations of line y=-8/3x+9 passes which through (-8,5)

OpenStudy (anonymous):

Parallel: lines will have the same slope Perpendicular: lines will have slopes that are inverse reciprocals of one another Parallel will be of form: y = -8/3x + b Solving for b using point given: 5 = -8(-8)/3 + b 5 = 64/3 + b 15/3 = 64/3 + b b = -49/3 Equation of the line is: y = -8/3 x - 49/3 Perpendicular will be of form: y = 3/8x + b 5 = -8(3)/8 + b 5 = -24/8 + b 5 = -3 + b b = 7 Equation of the line is: y = 3/8x + 7

OpenStudy (anonymous):

apparently the perpendicular one is incorrect

OpenStudy (anonymous):

the b was 8

OpenStudy (anonymous):

Yea, my apologies, you can see my poor algebra thinking 5+3 = 7 xD